# Learning Large-Scale Competitive Team Behaviors with Mean-Field Interactions and Online Opponent Modeling

### v2

@@ Appendix B Proof of Theoretical Results
For inequality (i), we used the inductive hypothesis; for inequality (ii), we utilized the Lipschitz continuity of the coordinator value function [Guan et al., 2024b, ] using mean-field distributions $(\mu^{\phi^{*}}_{t+1},\nu^{\psi^{\bar{N}_{2}}_{t}}_{\mathrm{apprx},t+1})$ from (B) and the the deterministic transitions from $t$ to $t+1$; inequality (iv) follows from the definition of the $\min$ operator; and equality (v) follows from the definition of $J_{\infty,t}^{\phi^{*},\tilde{\psi}^{*}_{\infty}}(\mu^{N_{1}}_{t},\nu^{N_{2}}_{t})$ which completes the induction.
@@
where ${J}_{\infty}^{\phi^{*}_{\infty},\psi^{*}_{\infty}}(\mu_{0},\nu_{0})=\max_{\phi\in\Phi}\min_{\psi\in\Psi}~J_{\infty}^{\phi,\psi}(\mu_{0},\nu_{0})$. Note that (i) follows from Step 2 (B.26); (ii) is a consequence of (B) and inequalities (iii) and (iv) follow from Lemmas 5 (applied on game $\mathcal{G}_{1}$) and 6 (applied on game $\mathcal{G}_{2}$) in Guan et al., 2024a respectively, combined with the fact that $\min(\bar{N}_{1},\bar{N}_{2})>\min(N_{1},N_{2})$, completing the proof. ∎
@@
where $\mathcal{M}_{t+1}=\mathcal{M}^{N_{1}}_{t}F(\mathcal{M}^{N_{1}}_{t},\mathcal{N}^{N_{2}}_{t},\phi^{*})$ and $\hat{\mathcal{M}}_{t+1}=\hat{\mathcal{M}}^{N_{1}}_{t}F(\hat{\mathcal{M}}^{N_{1}}_{t},\hat{\mathcal{N}}^{N_{2}}_{t},\phi^{*},\Gamma_{\textrm{D-PC}})$ are shorthand notations for the next induced mean-field from the infinite-population deterministic dynamics [Guan et al., 2024b, ] with and without the estimator $\Gamma_{\textrm{D-PC}}$. From Corollary 1 in Guan et al., 2024b , we have
@@
The exponential convergence of D-PC at each time-step $t$ follows directly from Theorems 6-8 in Nedić and Liu, [2016] as the number of communication rounds $R_{\textrm{com}}\to\infty$. Therefore, for finite communication rounds, there exist constants $c_{x},c_{y}>0$ and $\rho_{x},\rho_{y}\in(0,1)$ such that for any number of communication rounds $R_{\mathrm{com}}$ and any underlying distributions $(\hat{\mu}_{t}^{N_{1}},\hat{\nu}_{t}^{N_{2}})$ D-PC satisfies
@@
Definition 6 (Set Regularity [Nedić and Liu,, 2016]).
@@
Refer to Proposition 1 from Nedić and Liu, [2016]. ∎
Following Theorem 8 from Nedić and Liu, [2016], we have
@@ C.4 Reward Structure
Thus, by the generalized version of von Neumann’s minimax theorem Sorin, [2002], we conclude that the game value is unique, proving the second part of the proposition  [^2] Note: The optimal infinite horizon trajectory itself need not be unique (we have shown that $x_{1}$ can take a range of values).. ∎
@@ C.5 Implementation Details and Hyperparameters
The state distributions are represented as arrays that are concatenated together to form the global observation. This becomes the input to the critic network which consists of a single hidden layer of 64 neurons and two tanh activation functions. The output is a single value that is equal to the estimated value function. On the other hand, the actor-network consists of a single MLP layer of 64 neurons that is concatenated with the local agent observation. Additionally, the logits are converted to a probability distribution through a softmax layer. The dimension scales with $|\mathcal{A}|$. Both the actor and critic networks are initialized using orthogonal initialization [Huang et al.,, 2022].
The single-stage RPS game is trained for 5,000 time steps with the actor and critic learning rates set to 0.0005 and 0.001, respectively, which remain constant throughout training. The networks are updated using the ADAM optimizer Kingma and Ba, [2017] every 50 time steps for 10 epochs and a PPO clip value of 0.1. The entropy is decayed from 0.01 to 0.001 geometrically. We use an episode length of 1 after which the rewards are bootstrapped.
Moreover, since we have a single “team” buffer and the input/output dimensions are small, we do not use a mini-batch based update. For cRPS we use an episode length of 10 after which the rewards are bootstrapped. cRPS is trained using 200,000 time steps (=20,000 episodes) and is updated every 100 time steps. The algorithm was trained on a single NVIDIA GeForce RTX 3070 GPU and the training times are given in Tables 1 and 2.
@@ D.2 Interaction Between Agents
The transitions between states for agents belonging to both teams are characterized by their dynamics. These dynamics are probabilistic and depend on interactions among agents and are weakly coupled through their mean-field distributions. The weak coupling dynamics is keeping in line with the assumption in Guan et al., 2024b .
@@ D.4 Implementation and Hyperparameters
A fully connected (dense) layer takes the flattened output of the convolutional block and reduces it to 512 units. We have a single hidden layer that reduces the dimension further to 128 and then the output logits. The layers are separated by the tanh activation functions. Finally, the logits are converted to a probability distribution through a softmax layer. Both the actor and critic networks are initialized using orthogonal initialization [Huang et al.,, 2022]. The architectures of the shared-team actor and minimally-informed critic networks for this example are shown in Figures D.2 and D.3 respectively.
@@
All maps are trained using a single NVIDIA GeForce RTX 3070 GPU. The actor and critic learning rates are set to 0.0005 and 0.001 and both decay geometrically by a factor of 0.999. The networks are updated using the ADAM optimizer [Kingma and Ba,, 2017] with two mini-batches for 10 epochs and a PPO clip value of 0.1. The entropy coefficient is initialized to 0.01 and decays with a factor of 0.995.
@@ E.1 Rock-Paper-Scissors (RPS)
We first extend the two-player Rock-Paper-Scissors (RPS) game to a game played between two populations as described in Appendix C. The Nash equilibrium for this population-based RPS game is the uniform population distribution $[1/3,1/3,1/3]$ over the 3 states [Raghavan,, 1994].
We compare MF-MAPPO with DDPG-MFTG [Shao et al., 2024b, ] based on the training time, average test rewards and attainment of the computed Nash distributions for $N_{1}=N_{2}=1,000$ agents. We exclude MADDPG [Lowe et al.,, 2017] from our comparison, as it scales poorly to hundreds or thousands of agents due to its reliance on all agents’ local and global observations and actions as inputs to its critic networks. We include the training curve and performance for cRPS in Figure E.4 and Table 2 respectively, for reference.
From the learning curves in Figure E.4 one can see that the DDPG-MFTG algorithm failed to converge to the analytical game value of zero, while MF-MAPPO almost immediately attained the Nash game value. This corroborates with the results presented for cRPS in the main text. However, as shown in Table 1, MF-MAPPO does take slightly longer to train since, unlike DDPG-MFTG, since MF-MAPPO avoids mini-batch training, following Yu et al., [2022].
@@
Figure E.4: Training curves for RPS and cRPS.
@@
Table 2: Performance comparison for cRPS
Approach Training Time Average Reward NE Attained?
MF-MAPPO 2h 17min 15s -0.331
DDPG-MFTG 60h 49min 41s 3.774
@@ E.2.1 Validation Cases for MF-MAPPO
In the second scenario (Figure E.8), with the Red team evenly distributed at the corners, 30% of the Blue agents start at cell [2, 2] and 70% at cell [3, 1]. The Red team’s numerical advantage at [3, 3] forces the Blue agents to move around and regroup (Figures E.8(b)-E.8(f)). Once united, the Blue team’s numerical advantage forces the Red subgroup at [0,1] to disperse to avoid deactivation, allowing Blue to reach the target.
@@ E.3 Mean-Field Estimation for Grid World Navigation Using D-PC
We use an edgeless visibility graph and a subgrid-based communication graph. In the subgrid communication topology, an agent located at state $x$ can exchange information with its neighboring states arranged in a $3\times 3$ grid centered around $x$, as depicted in Figure E.13. Naturally, agents at the boundary have reduced connectivity: edge cells communicate within a $2\times 3$ neighborhood, and corner cells within a $2\times 2$ configuration, also shown in Figure E.13. We evaluate the estimation performance for communication rounds $R_{\mathrm{com}}\in\{1,2,3,4,5,6,7,8\}$ and consider population sizes of $N=100$ and $N=1,000$ (Recall, Theorem 3). From Figures E.14 and E.16 we see that up to approximately $R_{\mathrm{com}}=4$, the proposed algorithm consistently outperforms the benchmark, particularly in the early phases where communication is limited.
@@
The benchmark algorithm [Benjamin and Abate, 2025b, ] assumes that each state directly acquires the exact distributions of its communication neighbors $\mathcal{N}^{\mathrm{com}}(x)$. Consequently, it achieves better accuracy when $R_{\mathrm{com}}$ approaches $\mathrm{diam}(\mathcal{G}^{\mathrm{com}}_{t})$ communication rounds. While our method converges slightly more gradually, it offers notable practical advantages: it avoids relying on direct access to neighbor distributions—thereby preserving privacy—and performs competitively even under tighter communication constraints. Furthermore, for $R_{\mathrm{com}}<\mathrm{diam}(\mathcal{G}^{\mathrm{com}}_{t})$, the benchmark estimates unobserved states using a uniform distribution assumption—potentially leading to inaccurate representations of the actual mean-field. This inaccuracy is especially visible for small communication budgets ($R_{\mathrm{com}}=1,2,3$). Although increasing rounds mitigates this error, in real-world scenarios where fast decision-making is critical such as in real-time opponent modeling, large-scale communication may be infeasible. In contrast, our algorithm remains tractable, communication-efficient, and robust across a range of practical deployment settings, offering a compelling trade-off between estimation quality and operational cost. The reward plots in Figures E.15 and  E.17, corresponding to both population sizes, demonstrate that our method exhibits notably lower regret compared to the benchmark when evaluated against the fully observable policy. This improvement is particularly evident when $R_{\mathrm{com}}<\frac{1}{2}\mathrm{diam}(\mathcal{G}^{\mathrm{com}}_{t})$. For higher number of communication rounds our performance remains competitive with the benchmark. The plots also show that the estimation error incurred reduces with an increase in population size, predominantly noticeable at the peak around $t=100$. Furthermore, the increase in population size from $N=100$ to $N=1,000$ results in reduced noise in the total variation errors and smoother error plots, thereby resulting in smaller variance. Both these results corroborate with the finite population mean-field approximation guarantees presented from Theorem 5. Empirically, we assert that our algorithm is particularly well-suited for scenarios with limited communication budgets, especially when $R_{\mathrm{com}}<\frac{1}{2}\mathrm{diam}(\mathcal{G}^{\mathrm{com}}_{t})$, where it consistently delivers competitive accuracy with substantially reduced overhead.
@@ E.4 Non-Lipschitz Policies and Estimation
Navigation. We consider the navigation problem defined above and now focus on the cumulative total variation error, i.e., $\sum_{t}\mathrm{d}_{\mathrm{TV}}\big(\mu_{t},\hat{\mu}_{t}\big)$. Figure E.19 show the performance of our proposed D-PC algorithm and the benchmark method from Benjamin and Abate, 2025b . In both estimation settings, once again, gradient-regularization yields trajectories that closely resemble the fully-observable scenario (lower errors).

### v3

@@ Appendix B Proof of Theoretical Results
For inequality (i), we used the inductive hypothesis; for inequality (ii), we utilized the Lipschitz continuity of the coordinator value function [Guan et al., 2024b] using mean-field distributions $(\mu^{\phi^{*}}_{t+1},\nu^{\psi^{\bar{N}_{2}}_{t}}_{\mathrm{apprx},t+1})$ from (B) and the the deterministic transitions from $t$ to $t+1$; inequality (iv) follows from the definition of the $\min$ operator; and equality (v) follows from the definition of $J_{\infty,t}^{\phi^{*},\tilde{\psi}^{*}_{\infty}}(\mu^{N_{1}}_{t},\nu^{N_{2}}_{t})$ which completes the induction.
@@
where ${J}_{\infty}^{\phi^{*}_{\infty},\psi^{*}_{\infty}}(\mu_{0},\nu_{0})=\max_{\phi\in\Phi}\min_{\psi\in\Psi}~J_{\infty}^{\phi,\psi}(\mu_{0},\nu_{0})$. Note that (i) follows from Step 2 (B.26); (ii) is a consequence of (B) and inequalities (iii) and (iv) follow from Lemmas 5 (applied on game $\mathcal{G}_{1}$) and 6 (applied on game $\mathcal{G}_{2}$) in Guan et al. [2024a] respectively, combined with the fact that $\min(\bar{N}_{1},\bar{N}_{2})>\min(N_{1},N_{2})$, completing the proof. ∎
@@
where $\mathcal{M}_{t+1}=\mathcal{M}^{N_{1}}_{t}F(\mathcal{M}^{N_{1}}_{t},\mathcal{N}^{N_{2}}_{t},\phi^{*})$ and $\hat{\mathcal{M}}_{t+1}=\hat{\mathcal{M}}^{N_{1}}_{t}F(\hat{\mathcal{M}}^{N_{1}}_{t},\hat{\mathcal{N}}^{N_{2}}_{t},\phi^{*},\Gamma_{\textrm{D-PC}})$ are shorthand notations for the next induced mean-field from the infinite-population deterministic dynamics [Guan et al., 2024b] with and without the estimator $\Gamma_{\textrm{D-PC}}$. From Corollary 1 in Guan et al. [2024b], we have
@@
The exponential convergence of D-PC at each time-step $t$ follows directly from Theorems 6-8 in Nedić and Liu [2016] as the number of communication rounds $R_{\textrm{com}}\to\infty$. Therefore, for finite communication rounds, there exist constants $c_{x},c_{y}>0$ and $\rho_{x},\rho_{y}\in(0,1)$ such that for any number of communication rounds $R_{\mathrm{com}}$ and any underlying distributions $(\hat{\mu}_{t}^{N_{1}},\hat{\nu}_{t}^{N_{2}})$ D-PC satisfies
@@
Definition 6 (Set Regularity [Nedić and Liu, 2016]).
@@
Refer to Proposition 1 from Nedić and Liu [2016]. ∎
Following Theorem 8 from Nedić and Liu [2016], we have
@@ C.4 Reward Structure
Thus, by the generalized version of von Neumann’s minimax theorem Sorin [2002], we conclude that the game value is unique, proving the second part of the proposition  [^2] Note: The optimal infinite horizon trajectory itself need not be unique (we have shown that $x_{1}$ can take a range of values).. ∎
@@ C.5 Implementation Details and Hyperparameters
The state distributions are represented as arrays that are concatenated together to form the global observation. This becomes the input to the critic network which consists of a single hidden layer of 64 neurons and two tanh activation functions. The output is a single value that is equal to the estimated value function. On the other hand, the actor-network consists of a single MLP layer of 64 neurons that is concatenated with the local agent observation. Additionally, the logits are converted to a probability distribution through a softmax layer. The dimension scales with $|\mathcal{A}|$. Both the actor and critic networks are initialized using orthogonal initialization [Huang et al., 2022].
The single-stage RPS game is trained for 5,000 time steps with the actor and critic learning rates set to 0.0005 and 0.001, respectively, which remain constant throughout training. The networks are updated using the ADAM optimizer Kingma and Ba [2017] every 50 time steps for 10 epochs and a PPO clip value of 0.1. The entropy is decayed from 0.01 to 0.001 geometrically. We use an episode length of 1 after which the rewards are bootstrapped.
Moreover, since we have a single “team” buffer and the input/output dimensions are small, we do not use a mini-batch based update. For cRPS we use an episode length of 10 after which the rewards are bootstrapped. cRPS is trained using 200,000 time steps (=20,000 episodes) and is updated every 100 time steps. The algorithm was trained on a single NVIDIA GeForce RTX 3070 GPU and the training times are given in Tables 3 and 1.
@@ D.2 Interaction Between Agents
The transitions between states for agents belonging to both teams are characterized by their dynamics. These dynamics are probabilistic and depend on interactions among agents and are weakly coupled through their mean-field distributions. The weak coupling dynamics is keeping in line with the assumption in Guan et al. [2024b].
@@ D.4 Implementation and Hyperparameters
A fully connected (dense) layer takes the flattened output of the convolutional block and reduces it to 512 units. We have a single hidden layer that reduces the dimension further to 128 and then the output logits. The layers are separated by the tanh activation functions. Finally, the logits are converted to a probability distribution through a softmax layer. Both the actor and critic networks are initialized using orthogonal initialization [Huang et al., 2022]. The architectures of the shared-team actor and minimally-informed critic networks for this example are shown in Figures D.2 and D.3 respectively.
@@
All maps are trained using a single NVIDIA GeForce RTX 3070 GPU. The actor and critic learning rates are set to 0.0005 and 0.001 and both decay geometrically by a factor of 0.999. The networks are updated using the ADAM optimizer [Kingma and Ba, 2017] with two mini-batches for 10 epochs and a PPO clip value of 0.1. The entropy coefficient is initialized to 0.01 and decays with a factor of 0.995.
@@ E.1 Rock-Paper-Scissors (RPS)
We first extend the two-player Rock-Paper-Scissors (RPS) game to a game played between two populations as described in Appendix C. The Nash equilibrium for this population-based RPS game is the uniform population distribution $[1/3,1/3,1/3]$ over the 3 states [Raghavan, 1994].
We compare MF-MAPPO with DDPG-MFTG [Shao et al., 2024b] based on the training time, average test rewards and attainment of the computed Nash distributions for $N_{1}=N_{2}=1,000$ agents. We exclude MADDPG [Lowe et al., 2017] from our comparison, as it scales poorly to hundreds or thousands of agents due to its reliance on all agents’ local and global observations and actions as inputs to its critic networks. We include the training curve for cRPS in Figure E.5 for reference.
From the learning curves in Figures E.4 and E.5 one can see that the DDPG-MFTG algorithm failed to converge to the analytical game value of zero, while MF-MAPPO almost immediately attained the Nash game value. This corroborates with the results presented for cRPS (Table 1) in the main text. However, as shown in Table 3, MF-MAPPO does take slightly longer to train since, unlike DDPG-MFTG, since MF-MAPPO avoids mini-batch training, following Yu et al. [2022].
@@
Figure E.4: Training curve for RPS.
Figure E.5: Training curve for cRPS.
@@ E.2.1 Validation Cases for MF-MAPPO
In the second scenario (Figure E.9), with the Red team evenly distributed at the corners, 30% of the Blue agents start at cell [2, 2] and 70% at cell [3, 1]. The Red team’s numerical advantage at [3, 3] forces the Blue agents to move around and regroup (Figures E.9(b)-E.9(f)). Once united, the Blue team’s numerical advantage forces the Red subgroup at [0,1] to disperse to avoid deactivation, allowing Blue to reach the target.
@@ E.2.2 Comparison of MF-MAPPO with Baseline
While Table 2 reports the performance of the learned policies when evaluated against each other, Table 4 provides a complementary analysis of the sample efficiency of the baselines in the Battlefield environment. We define sample efficiency as the reward obtained per interaction with the environment:
$\textrm{Sample Efficiency}=\frac{\textrm{Total rewards}}{\textrm{Total environment steps}}.$
In addition, we report the average episode length, which serves as an indicator of how quickly episodes terminate. Shorter episodes suggest that the Blue agents increasingly learn to reach the target, while the Red agents simultaneously learn to defend the target and deactivate Blue agents—either outcome leading to episode termination.
Table 4: Performance metrics for Battlefield
Algorithm Critic Input Sample Efficiency Avg. Episode Length
MF-IPPO $\{(x^{N_{1}}_{i,t},\mu_{t})\}_{i=1}^{N_{1}}$ 6.3392 9.258
MAPPO-PS $\{(x^{N_{1}}_{i,t},\mu_{t},\nu_{t})\}_{i=1}^{N_{1}}$ 5.4581 12.084
MAPPO-$C$ $(\mathbf{x}^{N_{1}}_{t},\mathbf{y}^{N_{2}}_{t},\mu_{t},\nu_{t})$ 5.7878 9.8808
DDPG-MFTG $(\phi_{t},\mu_{t},\nu_{t})$ 0.0548 20.0
MF-MAPPO $(\mu_{t},\nu_{t})$ 14.6238 4.8448
@@ E.3 Mean-Field Estimation for Grid World Navigation Using D-PC
We use an edgeless visibility graph and a subgrid-based communication graph. In the subgrid communication topology, an agent located at state $x$ can exchange information with its neighboring states arranged in a $3\times 3$ grid centered around $x$, as depicted in Figure E.14. Naturally, agents at the boundary have reduced connectivity: edge cells communicate within a $2\times 3$ neighborhood, and corner cells within a $2\times 2$ configuration, also shown in Figure E.14. We evaluate the estimation performance for communication rounds $R_{\mathrm{com}}\in\{1,2,3,4,5,6,7,8\}$ and consider population sizes of $N=100$ and $N=1,000$ (Recall, Theorem 3). From Figures E.15 and E.17 we see that up to approximately $R_{\mathrm{com}}=4$, the proposed algorithm consistently outperforms the benchmark, particularly in the early phases where communication is limited.
@@
The benchmark algorithm [Benjamin and Abate, 2025b] assumes that each state directly acquires the exact distributions of its communication neighbors $\mathcal{N}^{\mathrm{com}}(x)$. Consequently, it achieves better accuracy when $R_{\mathrm{com}}$ approaches $\mathrm{diam}(\mathcal{G}^{\mathrm{com}}_{t})$ communication rounds. While our method converges slightly more gradually, it offers notable practical advantages: it avoids relying on direct access to neighbor distributions—thereby preserving privacy—and performs competitively even under tighter communication constraints. Furthermore, for $R_{\mathrm{com}}<\mathrm{diam}(\mathcal{G}^{\mathrm{com}}_{t})$, the benchmark estimates unobserved states using a uniform distribution assumption—potentially leading to inaccurate representations of the actual mean-field. This inaccuracy is especially visible for small communication budgets ($R_{\mathrm{com}}=1,2,3$). Although increasing rounds mitigates this error, in real-world scenarios where fast decision-making is critical such as in real-time opponent modeling, large-scale communication may be infeasible. In contrast, our algorithm remains tractable, communication-efficient, and robust across a range of practical deployment settings, offering a compelling trade-off between estimation quality and operational cost. The reward plots in Figures E.16 and  E.18, corresponding to both population sizes, demonstrate that our method exhibits notably lower regret compared to the benchmark when evaluated against the fully observable policy. This improvement is particularly evident when $R_{\mathrm{com}}<\frac{1}{2}\mathrm{diam}(\mathcal{G}^{\mathrm{com}}_{t})$. For higher number of communication rounds our performance remains competitive with the benchmark. The plots also show that the estimation error incurred reduces with an increase in population size, predominantly noticeable at the peak around $t=100$. Furthermore, the increase in population size from $N=100$ to $N=1,000$ results in reduced noise in the total variation errors and smoother error plots, thereby resulting in smaller variance. Both these results corroborate with the finite population mean-field approximation guarantees presented from Theorem 5. Empirically, we assert that our algorithm is particularly well-suited for scenarios with limited communication budgets, especially when $R_{\mathrm{com}}<\frac{1}{2}\mathrm{diam}(\mathcal{G}^{\mathrm{com}}_{t})$, where it consistently delivers competitive accuracy with substantially reduced overhead.
@@ E.4 Non-Lipschitz Policies and Estimation
Navigation. We consider the navigation problem defined above and now focus on the cumulative total variation error, i.e., $\sum_{t}\mathrm{d}_{\mathrm{TV}}\big(\mu_{t},\hat{\mu}_{t}\big)$. Figure E.20 show the performance of our proposed D-PC algorithm and the benchmark method from Benjamin and Abate [2025b]. In both estimation settings, once again, gradient-regularization yields trajectories that closely resemble the fully-observable scenario (lower errors).